OpenStudy (anonymous):

\[\lim_{x \rightarrow 0} ( 1+x+f(x)/x) ) ^{1/n} = e^3\] then find \[\lim_{x \rightarrow 0} \frac{ f(x) }{ x^2}\]

OpenStudy (anonymous):

@hartnn @RadEn @Coolsector

OpenStudy (anonymous):

Try by taking log of both sides of equation.

OpenStudy (anonymous):

*tried...no result

OpenStudy (ikram002p):

mmm I thing there is no result sice limit of 1/x is not exist ... retype the qs again

OpenStudy (anonymous):

log (1+x+f(x)/x)=3x..solve for f(x).

OpenStudy (anonymous):

Using L'Hospital's rule you get that the log term must be of order 3x.

OpenStudy (ikram002p):

ok i think you well goes like this\[(\lim 1+x+\frac{ f(x) }{ x } )^\frac{ 1 }{ n }= e^3 \] \[\lim 1+x+\frac{ f(x) }{ x } = e^3n \] \[\lim \frac{ f(x) }{ x } = e^3n -1\]

OpenStudy (anonymous):

oh i thought the n was x..sorry for that :(

OpenStudy (ikram002p):

ya i thouth it x at first too

OpenStudy (anonymous):

lol...Sorry it is x

OpenStudy (anonymous):

not n...:(

OpenStudy (ikram002p):

huh ... mmm use the log as @quarkine showed

OpenStudy (ikram002p):

but not log take ln

OpenStudy (anonymous):

yep i meant ln too.in math i guess we almost never bother with base 10!

OpenStudy (anonymous):

@satellite73 @bahrom7893

OpenStudy (anonymous):

@Mertsj